# Jointly modelling the evolution of community structure and language in online extremist groups

### v1

@@ Abstract
Group interactions take place within a particular socio-temporal context, which should be taken into account when modelling communities. We propose a method for jointly modelling community structure and language over time, and apply it in the context of extremist anti-women online groups (collectively known as the manosphere). Our model derives temporally grounded embeddings for words and users, which evolve over the training window. We show that this approach outperforms prior models which lacked one of these components (i.e. not incorporating social structure, or using static word embeddings). Using these embeddings, we investigate the evolution of users and words within these communities in three ways: (i) we model a user as a sequence of embeddings and forecast their affinity groups beyond the training window, (ii) we illustrate how word evolution is useful in the context of temporal events, and (iii) we characterise the propensity for violent language within subgroups of the manosphere.
@@ 1 Introduction
Computational approaches to modelling extremist groups is a pressing current topic of research. Two aspects in particular are often considered when investigating these groups: the structure of the community (who talks to whom) and the language they use (what do they say). For example, Gialampoukidis et al. [15] and Berzinji et al. [5] use network analyses to identify key members of terrorist organisations and Ferrara et al. [14] model the probability of users interacting with extremist content and accounts using network- and time-based features. The use of targeted hate speech has been investigated to identify the promoters of various extremist ideologies [24, 1].
@@
The premise of this work is that language and social structure are mutually informative expressions of group belonging that should be modelled simultaneously. Moreover, such interactions take place in a particular temporal context; on the internet, language changes at rapid rates [23] and the community itself morphs and changes focus in response to current events [4, 20].
@@
To model the interaction of these dimensions, we propose a shared matrix factorisation model which jointly encodes linguistic and social evolution over time, yielding dynamic user and word embeddings. We evaluate this model in the context of online anti-women groups (collectively referred to as the manosphere) and illustrate the usefulness of the resulting embeddings in researching these communities.
@@
The proposed model is an amalgamation of two prior works: Appel et al. [2] developed a shared matrix factorisation approach which captures the temporal evolution of users, but takes word representations as being static; while Yao et al. [25] use a similar method to construct dynamic word representations, but do not incorporate social information. By combining these two approaches, we obtain representations of users and words that evolve over time and are informed by their social context. This enables us to reason about the temporal evolution of users and their language in a shared space; a promising step to advance computational modelling of extremist groups. That said, this method is not specific to extremist groups and can be applied within any community for which social and linguistic information is available over multiple timesteps.
Experimental results (Section 6) show that our model represents a statistically significant improvement in terms of reconstruction fidelity and cluster purity over prior models. Furthermore, we illustrate that the resulting user embeddings can be used to predict future behaviour of an individual more accurately (Section 6.2). We show the importance of dynamic, temporally-grounded word embeddings in Section 7. Finally, we present a novel characterisation of violent language in subgroups within this community which leverages the shared user and word embedding space (Section 8).
@@
To summarise, our contributions in this paper are focused in two primary areas:
  • We propose an extension to the models of Appel et al. [2] and Yao et al. [25] and present experimental results showing that this model outperforms the prior works.
  • We illustrate its usefulness in the context of modelling online extremist communities and present novel insights on a particular extremist group.
2 The manosphere
The manosphere is broadly defined as a collection of communities with a common interest in men’s issues, who are known to engage in highly toxic online behaviour as well as acts of real-world violence [20]. While the roots of this community may be traced back to the Men’s Liberation Movement in the 60s and 70s, it has seen a resurgence in recent years [20].
Social structure
Though often referred to as a whole, the manosphere in fact consists of distinct subgroups [3]. Men’s Rights Activists (MRA) aim to form a counterpoise to feminist advocacy, while Pickup Artists (PuA) and The Red Pill (TRP) focus on personal development. MGTOW (“Men Going Their Own Way”) promote social separation from women, and Incels (“Involuntarily Celibates”) adopt a nihilistic mindset that promotes self-harm or harm to others.
These differing perspectives can cause friction within the manosphere, such as the Incel community’s contempt for the efforts of PuAs and MGTOW’s criticism of MRAs’ reform efforts [3]. This type of division is often referred to as “splintering” within counterterrorism research, defined by Baele et al. [4] as a process whereby extremist ideologies tend to fragment over time into a range of sub-ideologies supported by rival factions, with minority splinter groups going towards increasing radicalism. The Incel subculture has been described as a more extreme splinter of the manosphere [20, 4].
In addition to the group-level evolution, individuals in groups also evolve in their level of dedication to the community. While the definition of radicalisation is widely disputed by scholars, it is generally agreed that it occurs through a gradual process over time [7, 10], highlighting the importance of modeling temporal effects when studying these communities.
Language
Lexical innovation has been shown to be strongly tied to social connectedness in online groups [8] and has been used as an indicator for radicalisation [13, 9]. Bogetić [6] states that the manosphere, and in particular Incels, are prolific lexical innovators, creating new terminology at more rapid rates than other extremist groups. Per illustration, the manosphere lexicon of Farrell et al. [12] includes more than 36 hate terms for women. The rapid evolution of language online, as well as the connection between social identity and language, again illustrates that these factors should be recognised as interconnected and dynamic.
3 Model definition
Our proposed model jointly decomposes a social adjacency matrix and a language content matrix over multiple timesteps. The model is illustrated in Figure 1. In this section, we discuss its components and compare it to other formulations.
@@ 3.1 Source matrices
where $\theta_{i}$ is the set of discussion threads user $i$ has engaged in. Intuitively, this formulation considers users to be similar if they interact with similar people. Similarly, given a vocabulary of size $d$, we construct a language content matrix $C\in\mathbb{R}^{m\times d}$ such that each element in $C$ represents the usage of a word $z$ by a user $i$, defined by its pointwise mutual information (PMI):
@@ 3.2 Matrix factorisation
When decomposed independently from $A$, the resulting matrix $U$ consists of an embedding for each of the $m$ users, which will be similar for users who often interact with the same set of context users. Similarly, when decomposed independently from $C$, the resulting matrix $U$ consists of an embedding for each of the $m$ users, which will be similar for users who often use the same uncommon words. To learn user embeddings that are informed by both linguistic and social behaviours, we jointly decompose $A$ and $C$ such that the user embeddings $U$ need to be able to reconstruct both source matrices.
Temporal evolution
To allow for temporal evolution over $T$ timesteps, we learn time-based latent factors $U_{t}$, $V_{t}$ and $W_{t}$ for $t\in\{1,...,T\}$ based on $A_{t}$ and $C_{t}$. To encourage alignment between embeddings over different timesteps, we follow Appel et al. [2] and Yao et al. [25] in adding a temporal smoothing term (shown in Eq. 3) that discourages large variations across timesteps.
Learning objective
@@
$\begin{split}J=\sum_{t=1}^{T}&\Big{(}||A_{t}-U_{t}V_{t}^{T}||^{2}+||C_{t}-U_{t}W_{t}^{T}||^{2}\\ &+||V_{t}||^{2}+||U_{t}||^{2}+||W_{t}||^{2}\\ &+||U_{t+1}-U_{t}||^{2}+||W_{t+1}-W_{t}||^{2}\\ &+||V_{t+1}-V_{t}||^{2}\Big{)},\end{split}$ (3)
where $\lambda_{1}$ and $\lambda_{2}$ are regularisation hyperparameters that suppress large weights and large inter-timestep embedding variation, respectively. We refer to this system as Cerberus, the three-headed dog from Greek mythology.
@@ 3.3 Comparison to other systems
In our evaluation, we compare Cerberus against the architectures of Appel et al. [2] and Yao et al. [25] to deconstruct $A$ and/or $C$ as they are defined in Section 3.1. As mentioned in Section 1, the former differs from ours in that it lacks dynamic word and context embeddings, whereas the latter lacks the adjacency source matrix.
In addition to the architecture differences, our formulations of the $A$ and $C$ matrices also differ subtly from the base systems.
Appel et al. [2] constructs both $A$ and $C$ based on counts rather than PMI. Using PMI instead potentially provides a more discriminative source matrix, since the rarity of words is incorporated. They further use static $V$ and $W$ matrices, but define a similar temporal evolution mechanism in $U$, yielding dynamic user embeddings but static word embeddings. We refer to this model (with source matrices as defined in Section 3.1) as StatCont (static content representation).
The original system of Yao et al. [25] performs a decomposition on a word-word matrix over time to yield dynamic word embeddings, and does not account for users or social structure. It uses the standard matrix factorisation formulation of word embeddings [16] whereby each word is represented according to how often it co-occurs within a small local window with a set of context words. Their premise is that words are similar if they often co-occur with the same words. By contrast, Cerberus uses a social definition of word meaning, and considers two words as being similar if they share similar usage patterns between users. In our evaluations, we use this definition of $C$ in the architecture of Yao et al. [25] to yield dynamic word and user representations. This model incorporates some social information indirectly since $C$ contains user representations and user similarity is captured through similar language usage, but it does not incorporate the thread structure (who is talking to whom). We refer to this model as NoAdj (no adjacency representation).
@@ 3.4 Further optimisations
We use the architecture of Appel et al. [2] as basis in this work, and incorporate four further optimisations beyond the original system:
@@
We follow Yao et al. [25] in relaxing the non-negativity constraint in the matrix factorisation, as the task does not inherently require non-negativity in the learned embeddings.
@@
Both the adjacency and content matrices are highly sparse, with a sparsity ratio exceeding 0.99. For both source matrices, the data is analagous to an implicit feedback setting, since a zero does not imply a negative score but only signifies a lack of interaction. For this reason, we downweigh the losses originating from empty elements in the source matrices. We use a scaling parameter $c_{0}=0.01$ to balance the losses originating from the observed and non-observed elements.
@@
Bias terms are used in matrix factorisation applications to capture global trends [19]. For example, some context users in the adjacency matrix may be highly active, meaning that their column might have high values for all users. Interacting with such a context user is not a very informative feature for an individual; as such, a bias term is used to abstract this type of information.
4 Data
@@
We use the Reddit portion of the manosphere dataset of Ribeiro et al. [20]. Any dataset of online conversations with user and thread IDs could be used to train this system; however, the subreddit structure is useful as it provides labels against which to evaluate the communities discovered by our system (more details in Section 5).
@@
The training data consists of posts across 50 subreddits belonging to 5 manosphere categories (annotated by Ribeiro et al. [20]): Incels, Men Going Their Own Way (MGTOW), Pick-up Artists (PUA), The Red Pill (TRP), Men’s Rights Activists (MRA). It also includes posts from subreddits about mental health (r/depression, r/socialanxiety and r/suicidewatch) and subreddits that are critical of the manosphere (r/exredpill, r/thebluepill).
@@
We are interested in short- to medium-term phenomena and behaviours of communities; as such, we use monthly training windows. We select a training period of 9 months (April-December 2018), consisting of 4,354,116 posts.
@@
As our focus is on evolution over time, we filter the dataset to include only users with interactions across three or more of the included timesteps, resulting in a set of 33,880 users. To construct the adjacency matrix $A$, we further filter the context users (the columns of $A$) to include only the top 10,000 users in terms of post frequency to reduce the computational expense of the matrix factorisation. To construct the content matrix $C$, we include words which are used by more than 20 users; a total of 44,679 words. The number of posts and users per subreddit are shown in Table 1.

### v2

# Jointly modelling the evolution of social structure and language in online communities

@@ Abstract
Group interactions take place within a particular socio-temporal context, which should be taken into account when modelling interactions in online communities. We propose a method for jointly modelling community structure and language over time. Our system produces dynamic word and user representations that can be used to cluster users, investigate thematic interests of groups, and predict group membership. We apply and evaluate our method in the context of a set of misogynistic extremist groups. Our results indicate that this approach outperforms prior models which lacked one of these components (i.e. not incorporating social structure, or using static word embeddings) when evaluated on clustering and embedding prediction tasks. Our method further enables novel types of analyses on online groups, including tracing their response to temporal events and quantifying their propensity for using violent language, which is of particular importance in the context of extremist groups.
@@ 1 Introduction
Online communities present a unique opportunity for studying the evolution of groups. Two aspects are often considered when researching online communities: the structure of the community (who talks to whom; e.g. Gialampoukidis et al. 2017) and the language they use (what do they say; e.g. Danescu-Niculescu-Mizil et al. 2013). While most existing works tend to explore these facets in isolation, the premise of our work is that language and social structure are mutually informative expressions of group belonging that should be modelled jointly. Moreover, such interactions take place in a particular temporal context. On the internet, language changes at rapid rates (Stewart and Eisenstein 2018) and the community itself morphs and changes focus in response to current events (Baele, Brace, and Ging 2023; Ribeiro et al. 2021).
@@
To model the interaction of these dimensions, we propose a shared matrix factorisation model which jointly encodes linguistic and social evolution over time, yielding dynamic user and word embeddings. We evaluate this model in the context of misogynistic extremist groups (collectively known as the manosphere) and illustrate the usefulness of the resulting embeddings in researching communities.
@@
The proposed model is an amalgamation of two prior works: Appel et al. (2019) developed a shared matrix factorisation approach which captures the temporal evolution of users, but modelled language as being static; while Yao et al. (2018) used a similar method to construct dynamic word representations, but did not incorporate social information. By combining these two approaches, we obtain representations of users and words that evolve over time and are informed by their social context. This enables us to reason about the temporal evolution of users and their language in a shared space; a promising step to advance computational modelling of extremist groups. That said, this method is not specific to extremist groups and can be applied within any community for which social and linguistic information is available over multiple timesteps.
Experimental results (Section 7) illustrate that our method improves over the abovementioned methods in several evaluation settings. Firstly, we cluster the dynamic user embeddings and find that our model obtains a statistically significant improvement in cluster purity when evaluated against subreddit labels. Secondly, we illustrate that the user embeddings can be used to predict future behaviour of an individual more accurately. In particular, we forecast embeddings and predict the subreddits with which a user will engage beyond the training window, finding that our model outperforms prior approaches in three different test formulations.
Our method further enables novel types of analyses and visualisations of online communities, individuals within these groups, and their evolving interests and language. We provide a qualitative illustration of the importance of dynamic, temporally-grounded word embeddings in Section 7.3 by investigating current affairs with a temporal dimension; e.g. the MeToo movement. Finally, we present a novel characterisation of violent language in subgroups within the manosphere, which leverages the shared user and word embedding space (Section 8). Our results show that Incel-related groups have the highest propensity for violent language among the manosphere subgroups, and that there is evidence of smaller, more violent groups, characteristic of the “splintering” phenomenon in extremist groups.
2 Modelling users and words over time
The goal of this work is to construct dynamic word and user embeddings in a shared space by accounting for three factors: language, social dynamics, and time. Such representations can form the basis for analysing and predicting the behaviour of individuals and groups, which is of particular importance within extremist communities. In this section, we discuss the relationships between each pair of factors as explored in prior work, and highlight their significance in the context of extremist groups.
2.1 Time and groups
Online communities are highly dynamic, with continuous developments at the micro level (through users entering and leaving the group) as well as the macro level (through evolving sub-community structures). Group polarisation theory (Myers and Lamm 1976; Sunstein 2002) provides one explanation for this phenomenon, hypothesising that discussions within homogeneous groups intensify extremist positions. This intensification forces results in the departure of moderate voices, creating smaller and increasingly radical communities. This can result in “splintering”, a term used in counterterrorism research for the process whereby extremist ideologies tend to fragment over time into a range of sub-ideologies supported by rival factions, with minority splinter groups going towards increasing radicalism (Baele, Brace, and Ging 2023).
Fracturing in groups can also be the result of external pressures, such as deplatforming. Di Martino et al. (2024) map nine mainstream and alt-tech sites and demonstrate that user migration driven by moderation pressure produces ideologically homogeneous “echo-platforms” characterised by a greater prevalence of unreliable content and a heightened ideological uniformity. Similarly, Vu, Hutchings, and Anderson (2024) study a deplatforming attempt on a doxxing platform, finding that it primarily affected casual users rather than core members, while also attracting new users who exhibit higher levels of toxicity at the time of joining.
At the micro level, individuals also evolve over time in their level of dedication to a community. While the definition of radicalisation is widely disputed by scholars, it is generally agreed that it occurs through a gradual process over time (Bowman-Grieve 2010; Della Porta and LaFree 2012).
2.2 Groups and language
Language has been shown to be strongly tied to social connectedness in online groups, as it is used to signal in-group status (Drake 1980). Social identity theory (Tajfel et al. 1979) provides a basis for this dynamic, positing that individuals derive self-concept from their group memberships, leading them to accentuate distinctions between their in-group and relevant out-groups. Contemporary research in sociolinguistics has observed these dynamics in the context of online communities. For example, Zhang et al. (2018) show that communities with more specialised and dynamic language are more likely to retain their users. Longitudinal tracking of user lifecycles further shows that users linguistically align themselves with a community as they become embedded in it (Danescu-Niculescu-Mizil et al. 2013). This linguistic alignment wanes before departure, making shifting language a dynamic proxy for changing identification with the group.
In the context of extremism, the use of group-specific neologisms has been proposed as a proxy measure for the amount of radicalisation influence an individual is under (e.g. Fernandez, Asif, and Alani 2018). De Kock and Hovy (2024) found that the use of group jargon is highly correlated with post volume and social connectedness in extremist groups, and that early adoption of group language is predictive of eventual levels of radicalisation indicators.
2.3 Language and time
The development of language over time is well-established within NLP research (Wang et al. 2020). The internet, having a strong social component and reduced impetus for formality, is a particularly fertile domain for creating new words (Stewart and Eisenstein 2018) and for imbuing existing words with new meanings (Mendelsohn et al. 2023). Bogetić (2023) states that the manosphere are prolific lexical innovators, creating new terminology at more rapid rates than other extremist groups. Per illustration, the manosphere lexicon of Farrell et al. (2019) includes more than 36 hateful neologisms for women. The rapid development of new language creates challenges in moderating toxic content online, since it can be difficult for outsiders to interpret (Mendelsohn et al. 2023).
2.4 Our contribution: combining all factors
Our proposed model, detailed in Section 3, learns dynamic user and word representations in the same space, informed by both linguistic and social context. To our knowledge, we are the first to model these factors jointly in the context of online communities. Comparing to systems that use static word representations or neglect social information, we obtain a statistically significant improvement in three evaluation settings, indicating that it is important to account for all three abovementioned interactions. Representing a community in this way allows for novel types of analyses of extremist groups, as explored in Sections 7.2, 7.3 and 8. Since this method relies on matrix factorisation, a well-established unsupervised modelling methodology, it is computationally efficient, interpretable and easily extendible.
3 Model definition
@@
Our proposed model jointly models social structure and language by decomposing a social adjacency matrix and a language content matrix over multiple timesteps. The model is illustrated in Figure 1. In this section, we discuss its components, optimisations, and comparisons with other formulations.
@@ 3.1 Source matrices
where $\theta_{i}$ is the set of discussion threads user $i$ has engaged in. Intuitively, this formulation considers users to be similar if they interact with similar people. Similarly, given a vocabulary of size $d$, we construct a language content matrix $C\in\mathbb{R}^{m\times d}$ such that each element in $C$ represents the usage of a word $z$ by a user $i$, defined by its pointwise mutual information (PMI; Jurafsky 2000):
@@ 3.2 Matrix factorisation
When decomposed independently based on $A$, the resulting matrix $U$ consists of an embedding for each of the $m$ users, which will be similar for users who often interact with the same set of context users. Similarly, when decomposed independently based on $C$, the resulting matrix $U$ consists of an embedding for each of the $m$ users, which will be similar for users who often use the same uncommon words. To learn user embeddings that are informed by both linguistic and social behaviours, we jointly decompose $A$ and $C$ such that the user embeddings $U$ need to be able to reconstruct both source matrices.
To allow for temporal evolution over $T$ timesteps, we learn time-based latent factors $U_{t}$, $V_{t}$ and $W_{t}$ for $t\in\{1,...,T\}$ based on $A_{t}$ and $C_{t}$. To encourage alignment between embeddings over different timesteps, we follow Appel et al. (2019) and Yao et al. (2018) in adding a temporal smoothing term (shown in Eq. 3) that discourages large variations across timesteps.
@@
$\begin{split}J&=\sum_{t=1}^{T}||A_{t}-U_{t}V_{t}^{T}||^{2}+||C_{t}-U_{t}W_{t}^{T}||^{2}\\ &+\sum_{t=1}^{T}\lambda_{1}(||V_{t}||^{2}+||U_{t}||^{2}+||W_{t}||^{2})\\ &+\sum_{t=1}^{T-1}\lambda_{2}(||U_{t+1}-U_{t}||^{2}+||W_{t+1}-W_{t}||^{2}\\ &\hskip 19.91684pt+||V_{t+1}-V_{t}||^{2}),\end{split}$ (3)
where $\lambda_{1}$ and $\lambda_{2}$ are regularisation hyperparameters that suppress large weights and large inter-timestep embedding variation, respectively. We refer to this system as Cerberus.
@@ 3.3 Comparison to other systems
In our evaluation, we compare Cerberus to the architectures of Yao et al. (2018) and Appel et al. (2019) to deconstruct $A$ and/or $C$ as they are defined in Section 3.1. As mentioned in Section 1, the former method differs from ours in that it lacks the adjacency source matrix, whereas the latter lacks dynamic word and context embeddings. In addition to the architecture differences, our formulations of the $A$ and $C$ matrices also differ subtly from the base systems.
The original system of Yao et al. (2018) performs a decomposition on a word-word matrix over time to yield dynamic word embeddings, and does not account for users or social structure. It uses the standard matrix factorisation formulation of word embeddings (Goldberg and Levy 2014) whereby each word is represented according to how often it co-occurs within a small local window with a set of context words. Their premise is that words are similar if they often co-occur with the same words. By contrast, Cerberus uses a social definition of word meaning, and considers two words as being similar if they share similar usage patterns between users. In our evaluations, we use this definition of $C$ in the architecture of Yao et al. (2018) to yield dynamic word and user representations. This model incorporates some social information indirectly since $C$ contains user representations and user similarity is captured through similar language usage, but it does not incorporate the thread structure (who is talking to whom). We refer to this model as NoAdj (no adjacency representation).
Appel et al. (2019) constructs both $A$ and $C$ based on counts rather than PMI. Using PMI instead potentially provides a more discriminative source matrix, since the rarity of words is incorporated. Notably, their system is evaluated in the context of researcher networks, using article keywords to construct the content matrix rather than conversation data, meaning that their data is less sparse. They further use static $V$ and $W$ matrices, but define a similar temporal evolution mechanism in $U$, yielding dynamic user embeddings but static word embeddings. We refer to this model (with source matrices as defined in Section 3.1) as StatCont (static content representation).
@@ 3.4 Further optimisations
We use the architecture of Appel et al. (2019) as basis in this work, and incorporate four further optimisations beyond the architecture changes:
@@
We follow Yao et al. (2018) in relaxing the non-negativity constraint in the matrix factorisation, as the task does not inherently require non-negativity in the learned embeddings.
@@
Both the adjacency and content matrices are highly sparse, with a sparsity ratio exceeding 0.99. For both source matrices, the data is analogous to an implicit feedback setting, since a zero does not imply a negative score but only signifies a lack of interaction. For this reason, we downweigh the losses originating from empty elements in the source matrices. We use a scaling parameter $c_{0}=0.01$ to balance the losses originating from the observed and non-observed elements.
@@
Bias terms are used in matrix factorisation applications to capture global trends (Koren, Bell, and Volinsky 2009). For example, some context users in the adjacency matrix may be highly active, meaning that their column might have high values for all users. Interacting with such a context user is not a very informative feature for an individual; as such, a bias term is used to abstract this type of information.
4 Case study: The manosphere
4.1 Background
The models in this study are trained on data from the manosphere, which is broadly defined as a collection of communities with a common interest in men’s issues, who are known to engage in highly toxic online behaviour as well as acts of real-world violence (Baele, Brace, and Coan 2021). These communities have complicated, evolving social structures (Ribeiro et al. 2021) along with highly specialised and dynamic vocabularies (Farrell et al. 2019), which make them well-suited to this study. The roots of the modern manosphere can be traced back to the Men’s Liberation Movement in the 60s and 70s. Since its inception, the movement has fragmented and reformed into several subgroups with related but distinct ideologies (Bachaud 2024), also referred to as categories in this work. For example, the Men’s Rights Activists (MRA) aim to form a counterpoise to feminist advocacy, while Pickup Artists (PuA) and The Red Pill (TRP) focus on personal development with an aim to seduce women. MGTOW (“Men Going Their Own Way”) promote social separation from women, and Incels (“Involuntarily Celibates”) adopt a nihilistic mindset that promotes self-harm or harm to others. These differing perspectives can cause friction and splintering within the manosphere, as exemplified in the Incels’ contempt for the efforts of PuAs to attract partners, and MGTOWs’ criticism of the reform efforts of the MRAs (Bachaud 2024).
4.2 Data
@@
We use the Reddit portion of the manosphere dataset of Ribeiro et al. (2021) for training. Any dataset of online conversations with user and thread IDs could be used to train this system; however, the subreddit structure is useful here as it provides labels against which to evaluate the communities discovered by our system (more details in Section 5). We expect that the system would generalise well to any platform that provides some mechanism for constructing a social graph (for example, a follower network).
@@
The training data consists of posts across 50 subreddits belonging to the 5 abovementioned manosphere categories (annotated by Ribeiro et al. (2021)): Incels, MGTOW, PuA, TRP and MRA. It also includes subreddits that are critical of the manosphere (r/exredpill, r/thebluepill) and subreddits about mental health (r/depression, r/socialanxiety and r/suicidewatch). We retain posts from these subreddits in our training data as they form a useful control mechanism, since they are included in the above dataset and therefore contain the data from the same time period using the same preprocessing, but are expected to be noticeably disparate from the manosphere groups. The number of posts and users per subreddit are shown in Table 1.
@@
We are interested in short- to medium-term phenomena and behaviours of communities; as such, we use monthly training windows. We select a training period of 9 months (April–December 2018), consisting of 4,354,116 posts.
@@
As our focus is on evolution over time, we filter the dataset to include only users with interactions across three or more of the included timesteps, resulting in a set of 33,880 users. To construct the adjacency matrix $A$, we further filter the context users (the columns of $A$) to include only the top 10,000 users in terms of post frequency to reduce the computational expense of the matrix factorisation. To construct the content matrix $C$, we include words that are used by more than 20 users; a total of 44,679 words. As a background corpus to calculate the PPMI of the content matrices, we use a dataset of over 40 million Reddit posts by Dziri et al. (2019).
4.3 Data limitations
The Ribeiro et al. (2021) dataset was selected because it provides a comprehensive selection of manosphere subcommunities on Reddit, a multi-community platform, which provides proxy labels for our evaluation. This does, however, introduce some limitations. We recognise that extremist communities also gather in platforms that are not readily accessible to outsiders; for example, Telegram and Discord. Given the closed-form nature of these groups, lacking a multi-community structure, it is possible that our method may not generalise well to those groups. However, these groups do sustain a significant presence on Reddit (and other public platforms like Twitter), which provides a recruitment channel for new members. Studying how users relationally evolve in a multicommunity network can provide important insights into the process of radicalisation and radicalisation pathways. Furthermore, these methods are not specific to the study of extremist groups, but can be useful for modelling the evolution of online communities more broadly.
The data used to train our models originates from 2018. While this dataset cannot be expected to capture the latest developments in the manosphere, it provides a comprehensive selection of manosphere subcommunities, and serves to illustrate the utility of our approach. The thematic distinctions between the 5 high-level manosphere groups are not expected to have evolved significantly in the given time period.